OpenStudy (anonymous):

Will medal! Algebra 2 help please

OpenStudy (mathmate):

hmm

OpenStudy (mathmate):

Is the middle number 15 or 16?

OpenStudy (anonymous):

sorry bout that I did the math wrong because the original equation is 4(x+2)^2-3 and you end up with 4x^2+16+11 when youre done right? and then you have to graph it

OpenStudy (mathmate):

I like better 4(x+2)^2-3 than the other one!

OpenStudy (anonymous):

it says to graph the equation... idk which one though

OpenStudy (mathmate):

It is easier to graph!

OpenStudy (mathmate):

It tells me right away that the graph is concave up, and has the vertex at -2,-3.

OpenStudy (anonymous):

oh yah it is graphing the function not the standard for of it

OpenStudy (anonymous):

sorry bout that haha but I still need help... im terrible with functions

OpenStudy (mathmate):

Well, let f(x)=4*(x+2)^2-3 compare with the form f(x)=a(x-h)^2+k we see that the vertex (h,k) is at (-2,-3). With me so far?

OpenStudy (anonymous):

yah so far... just making sure... the vertex is the middle point of the parabola right? like the point in the middle

OpenStudy (mathmate):

Right, through the line of symmetry. It is also the trough in this case. We also know that the leading coefficient (x^2) has a positive coefficient of 16, so it is concave up.

OpenStudy (anonymous):

ok so then what

OpenStudy (mathmate):

Now we need the zeroes of the function f(x), and the y-intercept. After that we're done. Are you able to do the two last steps?

OpenStudy (anonymous):

what? is there some way you can show me? im better if I can see it... sorry

OpenStudy (anonymous):

well actually... how do we find the zeros of the function and the y int?

OpenStudy (anonymous):

@mathmate ?

OpenStudy (mathmate):

Well, you would use the quadratic formula to solve for the zeroes.

OpenStudy (mathmate):

That is when f(x)=4*(x+2)^2-3=0 you need to find the values of x when f(x)=0. The expanded version of the function is f(x)=4*x^2+16*x+13=0 use the quadratic formula to find the values of x. a=4 b=16 c=13 The zeroes are not integers.

OpenStudy (anonymous):

sorry I was out of the house @mathmate

OpenStudy (anonymous):

ok so -16+ or minus the square root of 16^2-4(4)(13) all divided by 2(8)

OpenStudy (mathmate):

And the y-intercept?

OpenStudy (anonymous):

ok so the end I got -16+ or - the sqrt of 48/ 16

OpenStudy (mathmate):

I think it's all divided by 8 instead of 16.

OpenStudy (anonymous):

oh yah sorry I multiplied by 2 2 times

OpenStudy (anonymous):

which equation do I use to find the y int?

OpenStudy (mathmate):

The y intercept is equal to f(0), which means that it is the constant term of the function.

OpenStudy (anonymous):

yah I know that but I don't know which equation to use? the f(x)=4(x+2)-3???

OpenStudy (mathmate):

After that you're done. I will show you a sketch where to fit things in. Annotate these points with an x-y pair, such as (0,13).

OpenStudy (anonymous):

I don't know what equation to use for the y int... is it the 4x^2 +16x+11 or the 4(x+2)-3

OpenStudy (anonymous):

@mathmate if I know which equation to use for the y int then I will be able to do it but I don't know which one to use... I think its the 4x^2+16x+13 so the y int would be 13 right?

OpenStudy (anonymous):

@mathmate come on were almost done! :) I wana finish this :)

OpenStudy (anonymous):

@Phebe

OpenStudy (phebe):

im here nd where's the question

OpenStudy (anonymous):

graph f(x)=4*(x+2)^2-3

OpenStudy (anonymous):